OpenStudy (anonymous):

Simplify This Expression. 3^-2 _______ 2^4

OpenStudy (anonymous):

-3*-3 is 9 2*2*2*2=16 9/16 is the simplified answer

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

\[\frac{ 3^{-2} }{ 2^{4} }=\frac{ 1 }{ 3^{2}2^{4} }=\frac{ 1 }{ 9\times16 }=\frac{ 1 }{ 144 }\] To get rid of a negative exponent flip it to the other side of the fraction. No, acex is wrong

OpenStudy (anonymous):

3^(-2) does NOT equal -3 x -3

OpenStudy (anonymous):

ohh yess

OpenStudy (anonymous):

\[3^{-2}=\frac{ 1 }{ 3^{2} }= \frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

didnt have to be a d*ck about it